Solve the equation for y: 2x+3y=9
Answer by mananth(16946)

2x+3y=9
add -2x
-2x+2x+3y=9-2x
3y=9-2x
divide by 3
3y/3 = 9/3-2x/3
y = 3-2x/3
...
m.ananth@hotmail.ca